OpenStudy (anonymous):

can somone help me in trying to get the are between these curves. I have : y^2-4x=4 and 4x-y=16. I solved them for x (cause i was gonna integrate them with respect to y) and got: x=(-1)+(y^2/4) and x=(4)-(y/4). I am not sure is those are right??

OpenStudy (anonymous):

And then i had the integral from -4 to 5 of -1+y^2/4-4+y/4

OpenStudy (anonymous):

ok, so did you integrate it?

OpenStudy (anonymous):

i am having problem b/c of the fractions

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

integrate it term by term.

OpenStudy (anonymous):

but what about the fractions

OpenStudy (anonymous):

\[\int f(x) + g(x) dx = \int f(x)dx + \int g(x)dx\]

OpenStudy (anonymous):

you can pull constant multiples out of an integral

OpenStudy (anonymous):

\[\int kf(x)dx = k\int f(x)dx\]

OpenStudy (anonymous):

so when you get to the terms with fractions, just pull out the 1/4 before you evaluate the integral

OpenStudy (anonymous):

so i got -5x+y^3/12+y^2/8

OpenStudy (anonymous):

why is there an x in there?

OpenStudy (anonymous):

oh its -5y

OpenStudy (anonymous):

ok, now evaluate that for your limits.

OpenStudy (anonymous):

man, i am still not getting the answer

OpenStudy (anonymous):

what are you getting?

OpenStudy (anonymous):

i am getting -225/8

OpenStudy (anonymous):

the books say it should be 243/8

OpenStudy (anonymous):

Well it certainly shouldn't be negative..

OpenStudy (anonymous):

is this right : y^2-4x=4 : x=-1+y^2/4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about this4x-y=16: x=4-y/4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so would the f(y)-g(y) look like this : -5+y^2/4-y/4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think the line is your f(y) and the quadratic is g(y) since the line is 'above' the parabola.

OpenStudy (anonymous):

yeah your right, the right most curve is the f(y)

OpenStudy (anonymous):

-y/4+4 - (y^2/4 - 1) = -y/4 + 4 - y^2/4 + 1

OpenStudy (anonymous):

can we add the 4 and the 1

OpenStudy (anonymous):

?